OpenStudy (anonymous):

can someone please answer these 2 problems and show work.... kn+4f=9v, solve for n 8c+6j=5p, solve for c

OpenStudy (gw2011):

kn+4f=9v kn=9v-4f n=(9v-4f)/k 8c+6j=5p 8c=5p-6j c=(5p-6j)/8

OpenStudy (anonymous):

thank you very much

OpenStudy (gw2011):

You're welcome